5. (Please if you are not willing to answer the question completely, leave the question to...

5.

(Please if you are not willing to answer the question completely, leave the question to someone who will!)

a. How long will it take for $ 500 to amount to $ 700, if you invest in the 8% compound quarterly?

b. A debt of $ 600 that must be paid within three years and another of $ 800 payable in four years, will be settled by means of a payment within two years. If the interest rate is 8% compounded semiannually, how much is the payment?
Answer #1
a Present value = $500
Future value = $700
Interest rate componded quarterly = 8%
Quarters required :-
Future value   =    Present value* (1+r/100)
700 = 500*(1+8/100)
= 4.36 periods/quarters
:- 1 period = 1 quarter
b Debt value = 600
Repayment period = 2 years
Compound semi annually = 8%
So future value will be to be paid = 600*(1+8/100)(1+8/100)(1+8/100)(1+8/100)
payment = 816.2934
Debt value = 800
Repayment period = 2 years
Compound semi annually = 8%
So future value will be to be paid = 800*(1+8/100)(1+8/100)(1+8/100)(1+8/100)
payment = 1088.391
Total payments = 1904.685